In the 2015 AFC championship game played at Gillette Stadium, the New England Patriots were accused of deflating footballs. In order for the ball pressure to decrease naturally without intentionally removing some of the air, the temperature of the gas in the balls would have had to change. If the game time temperature was 9.5 °C, what temperature would the balls have to have been inflated at for the pressure of the balls to drop from 27.2 psi to 25.2 psi?
1
Expert's answer
2019-11-13T01:09:03-0500

P1/T1=P2/T2.

T2=P2xT1/P1=25.2x282.5/27.2=261.73 K